OpenStudy (anonymous):

If you are finding the line perpendicular to x + 3y = 5, what is the slope of the perpendicular line?

OpenStudy (ivettef365):

do you know how to find slope ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@ivettef365

OpenStudy (ivettef365):

convert the equation into y=mx + b where m = slope

OpenStudy (ivettef365):

you have x + 3y = 5 3y = -x + 5 y = -1/3x + 5/3 now you have slope = - 1/3

OpenStudy (ivettef365):

now to be perpendicular you need the opposite reciprocal

OpenStudy (ivettef365):

opposite of negative is positive reciprocal of 1/3 is 3/1 or 3 so the reciprocal is 3

OpenStudy (ivettef365):

sorry the perpendicular is 3

OpenStudy (anonymous):

ok. I get it now. Thanks for your help.

OpenStudy (ivettef365):

your welcome :)